James Cameron's stuntman is performing a stunt for an upcoming holiday movie romance. In the stunt, he has to jump off the top of a very high bridge and land in the water below. At a point during the fall, when he is still 58.0 m above the water, James Cameron's stuntman is falling at a speed of 39 m/s. His mass is 64 kg.

a. How much total energy does he have at this point?

b. What will his velocity be when he hits the water?


1
Expert's answer
2021-08-02T08:40:54-0400

Given:

"h=58\:\rm m"

"v_i=39\:\rm m/s"

"m=64\:\rm kg"

(a)

"E=mgh+\frac{mv_i^2}{2}=64*9.8*58+\frac{64*39^2}{2}=85000\:\rm J"

(b)

"mgh+\frac{mv_i^2}{2}=\frac{mv_f^2}{2}"

"v_f=\sqrt{v_i^2+2gh}=\sqrt{39^2+2*9.8*58}=52\:\rm m/s"